Circular No. 5869

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 SB AND 1993 SC Iwan P. Williams, Queen Mary and Westfield College, London; and

Alan Fitzsimmons and Donal O’Ceallaigh, Queen’s University, Belfast, report their detection of two faint asteroidal objects during Sept. 16-21 very slowly retrograding ∼ 1-2^{o} west of the vernal equinox point. The observations were made with the 2.5-m Isaac Newton telescope and and prime-focus CCD camera at La Palma and are given on MPEC 1993-S09 and 1993-S10, from which the following orbital data are extracted:

Direct circular orbital elements and ephemeris for 1993 SB:

Epoch = 1993 Sept. 10.0 TT  u = 2.^{o}95
Ω = 355.36 }2000.0
a  = 33.1491 AU  i = 2.28
1993TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Sept. 20 23^{h}53.^{m}92 − 0^{o}31.^{′}5 32.145 33.149 178.^{o}7 0.^{o}0 23.2
30 23 53.00 − 0 37.3 32.159 33.149 171.2 0.3 23.2
Oct. 10 23 52.11 − 0 43.0 32.203 33.149 161.1 0.6 23.2

Direct circular orbital elements and ephemeris for 1993 SC:

Epoch = 1993 Sept. 10.0 TT  u = 3.^{o}80
Ω = 354.96 }2000.0
a  = 34.4515 AU  i = 5.58
1993TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Sept. 20 23^{h}55.^{m}11 − 0^{o}06.^{′}6 33.448 34.451 178.^{o}3 0.^{o}1 22.3
30 23 54.22 − 0 12.1 33.460 34.451 171.6 0.2 22.4
Oct. 10 23 53.35 − 0 17.5 33.503 34.451 161.6 0.5 22.4

As for 1993 RO and RP, the observations can also be satisfied by parabolic orbits. With now six very faint objects detected at current distances beyond Neptune, and with many more discoveries presumably likely, there is rapidly developing a very severe problem of securing adequate astrometric follow-up, which is absolutely essential for any understanding of this exciting development in the outer solar system. Because of the innate indeterminacy of the orbit solutions, the problem is much more severe than that involving main-belt asteroids in the early nineteenth century, and all efforts should be made to ensure coverage of each new object on two nights each month well through evening quadrature.

1993 September 22 Brian G. Marsden